### v1

# ActAlign: Zero-Shot Fine-Grained Video Classification via Language-Guided Sequence Alignment

Abstract
We address the task of zero-shot fine-grained video classification, where no video examples or temporal annotations are available for unseen action classes. While contrastive vision–language models such as SigLIP demonstrate strong open-set recognition via mean-pooled image–text similarity, they fail to capture the temporal structure critical for distinguishing fine-grained activities. We introduce ActAlign, a zero-shot framework that formulates video classification as sequence alignment. For each class, a large language model generates an ordered sub-action sequence, which is aligned with video frames using Dynamic Time Warping (DTW) in a shared embedding space. Without any video–text supervision or fine-tuning, ActAlign achieves 30.5% accuracy on the extremely challenging ActionAtlas benchmark, where human accuracy is only 61.6%. ActAlign outperforms billion-parameter video–language models while using approximately 8$\times$ less parameters. These results demonstrate that structured language priors, combined with classical alignment techniques, offer a scalable and general approach to unlocking the open-set recognition potential of vision–language models for fine-grained video understanding.
Codehttps://github.com/aghdamamir/ActAlign
1 Introduction
Understanding fine-grained human activities in video—such as distinguishing a hook shot from a layup in basketball, or recognizing tactical formations in football—requires parsing subtle, temporally ordered visual cues across frames. These actions unfold in structured sequences of sub-events and are often nearly indistinguishable from one another in appearance. In contrast to general activities like swimming, which can often be inferred from a single frame showing a person in water, fine-grained recognition demands attention to temporally extended object interactions, spatial relations, and high-level intent. As such, models must not only understand what is present in a video but also when and how key sub-actions occur. This requires accurately aligning the temporal progression of sub-actions with each fine-grained activity to ensure a correct prediction, as shown in Figure 1.
At the same time, contrastive vision–language models such as CLIP (Radford et al. 2021) and SigLIP (Zhai et al. 2023) have demonstrated impressive open-set fine-grained recognition in static image domains by training on massive image–text pairs. These contrastive models enable zero-shot classification using natural language prompts and have been widely adopted for downstream recognition tasks. However, extending these capabilities to video understanding introduces new challenges and requires temporal modeling. Existing methods that adapt CLIP-style models to video recognition either average frame-level features (Rasheed et al. 2023; Zohra et al. 2025)—ignoring temporal structure—or fine-tune on target datasets (Wang, Xing, and Liu 2022; Ni et al. 2022; Kim et al. 2024; Wang et al. 2024a), sacrificing generalization and open-set recognition. In both cases, the fine-grained temporal semantics of actions are lost or diluted.
Recent video–language architectures and instruction-tuned LLM-based systems such as Video-LLaMA (Zhang, Li, and Bing 2023), VideoChat (KunChang Li 2023) mPLUG-Owl (Ye et al. 2023), Qwen2-VL (Wang et al. 2024b), and DeepSeek-JanusPro (Chen et al. 2025) enable open-ended, dialog-style video understanding through heavy instruction tuning, but they are not tailored for fine-grained video recognition.
Meanwhile, textual grounding (image–text alignment) remains a central challenge in interactive video-language models, especially for open-set and fine-grained video recognition. Dynamic Time Warping (DTW) (Vintsyuk 1968), a classical algorithm for aligning temporally mismatched sequences, has seen renewed interest through differentiable variants (Chang et al. 2019; Dogan et al. 2018) designed for supervised image–text temporal alignment. Yet, these methods rely on annotated transcripts or example support videos, making them impractical for zero-shot recognition. Likewise, approaches using part-level or attribute-level supervision (Wu et al. 2023; Zhu et al. 2024) offer fine-grained cues but lack the ability to model the temporal structure between language-defined actions and visual content.
Image-Text Embedding Space
Figure 1: ActAlign significantly improves zero-shot fine-grained action recognition by modeling them as structured language sequences. By aligning sub-action descriptions with video frames (green vs. red paths), we achieve more accurate predictions without requiring any video-text training data.
Figure 2: Our pipeline uses an LLM to generate structured sub-action sequences for each fine-grained candidate class in ActionAtlas (Salehi et al. 2024), forming SubActionAtlas. This structured representation enables alignment with video content for zero-shot recognition tasks.
In this work, we introduce ActAlign, a novel framework that brings the open-set generalization power of image–text models to fine-grained video recognition through language-guided subaction alignment in a truly zero-shot setting. Rather than tuning a model for a specific domain or collapsing the video into a static representation, ActAlign operates in a training-free setting: for each unseen action class, we use a large language model (LLM) to generate a structured sequence of temporal sub-actions that semantically define the class. (see Figure 2. Then, using the pretrained SigLIP model (Zhai et al. 2023) to extract frame-wise visual and subaction features, we align the video sequence with the LLM-generated subaction script via Dynamic Time Warping (DTW).(see Figure 1) This allows us to compute a soft alignment score between different action classes that respects both content and temporal ordering, enabling fine-grained classification in a truly zero-shot manner.
To rigorously evaluate our approach, we use ActionAtlas (Salehi et al. 2024), the most challenging fine-grained video recognition benchmark to our knowledge, with a human accuracy ceiling of 61.64%. It comprises sports footage paired with extremely fine-grained candidate tactics that demand close attention to subtle visual cues for correct classification. We construct SubActionAtlas, by decomposing each candidate action into a sequence of sub-actions via LLM prompting (Figure 2), providing structured templates for alignment.
Our contributions are as follows:
  • We introduce a novel framework for fine-grained video recognition that models each action as a general, structured temporal sequence of sub-actions derived solely from action names—without access to videos or transcripts.
  • We propose ActAlign, a novel zero-shot framework that applies the open-set generalization strength of contrastive image–text models to the challenging task of fine-grained and open-set video classification, without requiring any video–text supervision.
  • We show that ActAlign consistently outperforms prior zero-shot and CLIP-based baselines, and even exceeds billion-parameter video–language models on challenging fine-grained video classification benchmarks.
2 Related Work
2.1 Contrastive Image–Language Models
Contrastive vision–language models such as CLIP (Radford et al. 2021), SigLip (Zhai et al. 2023), and ALIGN (Jia et al. 2021) learn joint image–text embeddings from large-scale image–caption pairs, enabling strong open-set recognition without task-specific training. These models serve as pretrained backbones for downstream tasks, including visual question answering and reasoning (Li, Baldridge, and Hoi 2021; Li et al. 2022; Tsimpoukelli et al. 2021), image captioning and generation (Mokady, Hertz, and Bermano 2021; Wang et al. 2021b), and few-shot and zero-shot classification (Zhou et al. 2021; Khattak et al. 2025). Notably, SigLip (Zhai et al. 2023) recently enhanced CLIP’s architecture and leveraged a larger training corpus to enhance in image–text matching (Zhai et al. 2023). However, such models lack any temporal structure for processing video inputs. This frame-level abstraction limits their open-set classification ability for video inputs.
2.2 Video–Language Models
Video-Language Modeling
Initial efforts to extend vision–language models to video focused on pairing visual frames with corresponding narrations or transcripts. MIL-NCE (Miech et al. 2020) introduced multi-instance contrastive learning to align videos and narrations in uncurated instructional datasets. Later approaches leveraged transformer architectures to model temporal sequences. VideoBERT (Sun et al. 2019) and ActBERT (Zhu, Xu, and Yang 2020) applied masked language modeling over sequences of video frames and transcripts, To improve computational efficiency, ClipBERT (Lei et al. 2021) proposed sparse frame sampling for end-to-end video–text alignment.
Interactive Video Language Models
Recent work integrates large language models (LLMs) with visual encoders to support open-ended video understanding tasks such as captioning and dialogue. Systems like Video-LLaMA (Zhang, Li, and Bing 2023), mPLUG-Owl (Ye et al. 2023), and VideoChat (KunChang Li 2023) combine pretrained visual backbones with chat-centric LLMs to generate detailed responses and spatiotemporal reasoning in conversational settings. For example, VideoChat enables multi-turn dialogue grounded in video content, while Video-LLaMA augments an LLM with temporal and audio query transformers. These systems excel at descriptive and interactive tasks but require extensive instruction tuning (Zhang, Li, and Bing 2023; KunChang Li 2023) and are not optimized for fine-grained video recognition.
CLIP for Video Classification
Pretrained CLIP-style Image-Language models exhibit strong open-set image recognition capability, motivating adaptations to utilize their power for video recognition tasks. These adaptations include temporal modules and special prompting strategies. ActionCLIP (Wang, Xing, and Liu 2022) introduced a “pre-train, prompt, and fine-tune” strategy, augmenting CLIP with handcrafted label prompts and training on action datasets. X-CLIP (Ni et al. 2022) extended CLIP with temporal attention layers to process frame sequences. Other adaptations inject motion-awareness via learnable visual prompts or adapters (Ju et al. 2022; Lin et al. 2022). While these approaches show strong performance on closed-set benchmarks like Kinetics, their effectiveness relies on dataset-specific fine-tuning. As shown in ViFi-CLIP (Rasheed et al. 2023), such specialization often harms zero-shot generalization by overfitting to seen classes.
2.3 Video-Text Sequence Alignment
Classic Alignment Methods
Aligning sequences of video frames to textual descriptions has long been a challenge, especially under weak supervision where frame-level labels are unavailable. Classic approaches often rely on Dynamic Time Warping (DTW) or similar sequence alignment algorithms. For example, Bojanowski et al. (Bojanowski et al. 2015) proposed aligning videos to ordered scripts by enforcing the temporal order of events, without precise timestamps. Other works adopt the Connectionist Temporal Classification (CTC) loss—originally developed for speech recognition—for action segmentation in video. Richard et al. (Richard, Kuehne, and Gall 2018) and Huang et al. (Huang et al. 2016) used CTC and Viterbi decoding to align video frames with a sequence of action labels, learning to segment actions without frame-level supervision. However, these approaches typically assume access to ground-truth transcripts (i.e., the precise ordered list of sub-actions) for every training video.
Dynamic Time Warping (DTW) for Video Classification
Beyond supervision regimes, alignment algorithms themselves have evolved to improve flexibility and learning capacity. Differentiable variants of DTW, such as D3TW (Chang et al. 2019), introduced continuous relaxations that enable end-to-end gradient-based optimization under ordering constraints. NeuMATCH (Dogan et al. 2018) proposed a neural alignment model using recurrent moving windows to align long video sequences to textual inputs. More recently, graph-based models such as the Dynamic Graph Warping Transformer (Wang et al. 2021a) have incorporated structured reasoning and global constraints to improve alignment accuracy. OTAM (Cao et al. 2020) utilizes DTW for few-shot video classification. These methods still rely on task-specific supervision such as paired video–text exemplars or support videos from the target classes.
Our work
Unlike prior work that either discards temporal structure, fine-tunes away generality, or assumes support data for alignment, ActAlign bridges vision and language via sequence-to-sequence matching—paving the way for generalizable, interpretable, and temporally grounded video understanding in zero-shot settings. Our method targets zero-shot video classification of fine-grained actions by leveraging language as a structured latent representation aligned to visual content, without requiring video-text training data or fine-tuning.
3 Method
3.1 Problem Definition
Let $\mathcal{D}={(V_{i},y_{i})}_{i=1}^{N}$ denote a dataset of $N$ videos, where each video $V_{i}$ is associated with a ground-truth fine-grained class label $y_{i}$ drawn from a set of $M$ candidate classes $\mathcal{Y}={c_{1},\dots,c_{M}}$. Each video $V_{i}$ consists of a sequence of $T_{i}$ frames as defined in Eq. 1:
$V_{i}=\{\mathbf{v}_{i}^{t}\}_{t=1}^{T_{i}},\quad\mathbf{v}_{i}^{t}\in\mathbb{R}^{H\times W\times 3},$ (1)
where $H$ and $W$ denote the frame height and width. In our zero-shot setting, no video examples of the target classes $\mathcal{Y}$ are used for training or tuning; only high-level action names $c_{j}$ are provided. The goal is to construct a function $f:\mathcal{V}\times\mathcal{Y}\to\mathbb{R}$ that effectively maps the sequence of video frames into their correct action class $y$. The predicted class label $\hat{y}_{i}$ for a video $V_{i}$ is given by Eq. 2:
$\hat{y}_{i}=\arg\max_{c_{m}\in\mathcal{Y}}f(V_{i},c_{m}).$ (2)
To leverage semantic priors from LLMs, we automatically decompose each class label $c_{m}$ into an ordered, variable-length sequence of $K_{m}$ textual sub-actions, as defined in Eq. 3:
$S_{m}=[s_{m,1},s_{m,2},\dots,s_{m,K_{m}}],$ (3)
where $s_{m,k}$ is a concise natural-language description of the $k$-th step in executing action $m$.
3.2 Our Method
We define $f(V_{i},c_{m})$ as the alignment score between the visual frame embeddings $\{\mathbf{v}_{i}^{t}\}$ and the sub-action sequence $S_{m}$, computed via Dynamic Time Warping (DTW). This alignment is performed in the image–text embedding space, without requiring any fine-tuning or video examples from the target label set.  Figure 3 illustrates the pipeline of our proposed approach.
Figure 3: Our ActAlign Method Overview. (1) Subaction Generation: Given fine-grained actions (e.g. Basketball Tactics), we prompt an LLM to decompose each action (e.g. Hookshot, JumpShot, Dunk) into a temporal sequence of sub-actions. (2) Temporal Alignment: Video frames are encoded by a frozen pretrained vision encoder and smoothed via a moving‐average filter. Simultaneously, each subaction is encoded by the text encoder. We compute a cosine‐similarity matrix between frame and subaction embeddings, then apply Dynamic Time Warping (DTW) to find the optimal alignment path and normalized alignment score. (3) Class Prediction: We repeat this process for each candidate action m, compare normalized alignment scores $\hat{\gamma}_{\text{video},m}$, and select the action sequence with the highest score as the final prediction.
3.3 Preliminary Subaction Generation by LLM
In domains requiring extremely fine-grained distinctions— for example, differentiating between tactical plays in sports footage—the high-level action class name $c_{m}$ alone often lack sufficient discriminatory power. To reveal subtle inter-class differences, it is critical to decompose each action into an ordered sequence of granular, trackable sub-actions. We therefore define a mapping $\mathcal{P}:\mathcal{Y}\to\mathcal{S}$, where $\mathcal{S}$ is the space of all possible sub-action scripts, and for each fine-grained action $c_{m}$, $\mathcal{P}(c_{m})=S_{m}=[s_{m,1},\dots,s_{m,K_{m}}]$ denotes the LLM-generated sequence of sub-actions.
To instantiate $\mathcal{P}$, we employ a pretrained large-language model (GPT-4o) via carefully engineered natural language prompts. Given the set of candidate action class names ${c_{1},\dots,c_{M}}$, our prompt instructs the model to:
  1. 1.
    Decompose it into a varriable-sized sequence of self-sufficient, observable steps (sub-actions) that are semantically coherent and temporally ordered.
  2. 2.
    Return the list $[s_{m,1},\dots,s_{m,K_{m}}]$ for each $c_{m}$ in a consistent and structured format.
We then parse the model’s output to form $S_{m}$. By leveraging the LLM’s extensive world knowledge and linguistic priors, this process yields high-quality subaction sequence without any manual annotation or video examples. These scripts serve as the semantic reference signals for subsequent temporal alignment and classification.
Importance of Context in Sub-Actions.
We find that terse sub-action descriptions (e.g., “drive forward”) are too ambiguous that can refer to several sports, making DTW alignment ineffective. In contrast, context-rich prompts (e.g., “drive forward to the rim in basketball”) yield semantically grounded sub-actions that align well with video frames and support effective sequence matching.
3.4 Visual and Semantic Feature Encoding
Once each class label $c_{m}$ is decomposed into its sub-action sequence $S_{m}$, we project both video frames and sub-actions into a shared $d$-dimensional embedding space using the pre-trained SigLIP image–text model, which is recognized for its strong zero-shot recognition performance.
Visual Embeddings
Let $\phi_{v}:\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{d}$ denote the vision encoder. For video $V_{i}=\{\mathbf{v}i^{t}\}_{t=1}^{T_{i}}$, we compute frame-level embeddings as shown in Eq. 4:
$\mathbf{z}_{i}^{t}=\phi_{v}(\mathbf{v}_{i}^{t})\in\mathbb{R}^{d},\quad t=1,\dots,T_{i}.$ (4)
Stacking yields $Z_{i}=[\mathbf{z}_{i}^{1},\dots,\mathbf{z}_{i}^{T_{i}}]\in\mathbb{R}^{d\times T_{i}}$.
Semantic Embeddings
Let $\phi_{t}:\mathcal{T}\to\mathbb{R}^{d}$ be the text encoder. For each class $c_{m}$ and its subaction sequence $S_{m}$, we embed each step as shown in Eq. 5:
$\mathbf{u}_{m,k}=\phi_{t}(s_{m,k})\in\mathbb{R}^{d},\quad k=1,\dots,K_{m},$ (5)
stitching into $U_{m}=[\mathbf{u}_{m,1},\dots,\mathbf{u}_{m,K_{m}}]\in\mathbb{R}^{d\times K_{m}}$.
Shared Latent Space
We normalize all embeddings so that the similarity
$\text{sim}(\mathbf{z}_{i}^{t},\mathbf{u}_{m,k})=\frac{\mathbf{z}_{i}^{t}\top\mathbf{u}_{m,k}}{|\mathbf{z}_{i}^{t}|\cdot|\mathbf{u}_{m,k}|}$ (6)
is a valid cosine similarity measure between frame $\mathbf{v}_{i}^{t}$ and sub-action $s_{m,k}$ of class $c_{m}$ (see Eq. 6). This cross-modality similarity forms the basis for alignment in the next step.
3.5 Dynamic Time Warping
After feature encoding, each video yields a visual embedding sequence $Z_{i}$ and ordered sequences of sub-action embeddings $U_{m}$ for each class $c_{m}$. We treat $U_{m}$ as the reference semantic signal and $\widetilde{Z}_{i}$ as the query visual signal. The $U_{m}$ could also be viewed as a prototype sequence for class $c_{m}$.
Signal Smoothing
Real-world footage often contains abrupt scene changes or irrelevant frames (e.g., replays, advertisements) that introduce noise into $Z_{i}$. To mitigate this, we apply a 1D moving-average filter of width $w$ across the temporal dimension, as defined in Eq. 7:
$\widetilde{\mathbf{z}}_{i}^{t}=\frac{1}{w}\sum_{\tau=t-\lfloor w/2\rfloor}^{t+\lfloor w/2\rfloor}\mathbf{z}_{i}^{\tau},$ (7)
with boundary conditions handled via zero padding. The kernel width $w$ controls the trade-off between noise reduction and temporal resolution.
Affinity Matrix Construction
Let the smoothed visual embeddings for video $V_{i}$ be $\widetilde{Z}_{i}=[\widetilde{\mathbf{z}}_{i}^{1},\dots,\widetilde{\mathbf{z}}_{i}^{T_{i}}]$ and the sub-action sequence embeddings for class $c_{m}$ be $U_{m}=[\mathbf{u}_{m,1},\dots,\mathbf{u}_{m,K_{m}}]$. We first compute the raw cosine similarity matrix as shown in Eq. 8:
$A^{(m,i)}_{k,t}=\bigl{\langle}\mathbf{u}_{m,k},\,\widetilde{\mathbf{z}}_{i}^{t}\bigr{\rangle},\quad A^{(m,i)}\in\mathbb{R}^{K_{m}\times T_{i}}$ (8)
[…]
$\hat{A}^{(m,i)}_{k,t}=\sigma\bigl{(}\alpha\,A^{(m,i)}_{k,t}+\beta\bigr{)},$ (9)
where $\alpha,\beta$ are learned scaling parameters. The resulting $\hat{A}^{(i,m)}$ is used as the input affinity matrix for DTW alignment.
DTW Alignment and Scoring
Given the Affinity matrix $\hat{A}^{(m,i)}$ (defined in Eq. 9), we seek a warping path $W^{(m,i)}=\{(k_{1},t_{1}),\dots,(k_{L},t_{L})\}$ that maximizes cumulative similarity under monotonicity and continuity constraints, as formulated in Eq. 10:
$W^{(m,i)}$ $=\arg\max_{W}\sum_{(k,t)\in W}\hat{A}^{(m,i)}_{k,t},$
s.t. W is a valid warping path between [$\textit{W is a valid warping path between }[1,K_{m}]\textit{ and }[1,T_{i}].$ (10)
We solve this using dynamic programming, as defined in Eq. 11:
$D_{k,t}=\hat{A}^{(m,i)}_{k,t}+\max\{D_{k,t-1},D_{k-1,t},D_{k-1,t-1}\},$ (11)
with the base case $D_{0,*}=D_{*,0}=-\infty$. The final alignment score is $\max_{k,t}D_{k,t}$, and backtracking recovers the optimal warping path $W^{(m,i)}$.
Prediction
Upon obtaining the optimal warping path $W^{(m,i)}$ between the smoothed visual sequence $\widetilde{Z}_{i}$ and the sub-action sequence $U_{m}$ for candidate class $c_{m}$, we compute the raw alignment score as defined in Eq. 12:
$\gamma_{i,m}=\sum_{(k,t)\in W^{(m,i)}}\hat{A}^{(i,m)}_{k,t},$ (12)
where $\hat{A}^{(i,m)}$ is the affinity matrix introduced in Eq. 9. To mitigate the bias toward longer warping paths (which can accumulate higher raw scores), we normalize $\gamma_{i,m}$ by the path length, resulting in the average alignment score (Eq. 13):
$\hat{\gamma}_{i,m}=\frac{1}{|W^{(i,m)}|}\gamma_{i,m},$ (13)
where $|W^{(i,m)}|$ is the number of matched frame–subaction pairs. Since the similarity values in $\hat{A}^{(i,m)}$ lie in $[0,1]$ (due to the sigmoid in Eq. 9), the normalized alignment score $\hat{\gamma}_{i,m}$ also lies in the range $[0,1]$.
Finally, we predict the class whose sub-action sequence best aligns—on average—with the observed video frames. This is done by selecting the class with the highest normalized alignment score (Eq. 14):
$\hat{y}_{i}=\arg\max_{c_{m}\in\mathcal{Y}}\;\hat{\gamma}_{i,m}.$ (14)
4 Experiment
Prompt Type Avg. Subactions Avg. Subactions / Domain Avg. Words / Subaction
Short-Fixed ($T=1$) $10.00\pm 0.00$ $10.00\pm 0.00$ $2\pm 0.03$
Context-Rich ($T=1$) $4.94\pm 0.86$ $5.01\pm 0.62$ $13.68\pm 2.78$
Context-Rich ($T=0.2$) $5.04\pm 0.87$ $5.09\pm 0.63$ $13.13\pm 2.55$
Table 1: Linguistic complexity of sub-action scripts generated by different prompting strategies. Context-rich prompts yield fewer but more descriptive sub-actions with significantly higher word counts compared to the Short-Fixed strategy. Lower temperature slightly increases consistency without sacrificing expressiveness.
Method #Param Top-1 (%) Top-2 (%) Top-3 (%)
Random (10 Trials) - $20.81$ $42.04$ $62.50$
Human Evaluation (Oracle) - $61.64$
mPLUG-Owl-Video (Ye et al. 2023) $7B$ $19.49$
VideoChat2 (KunChang Li 2023) $7B$ $21.27$
VideoLLaMA (Zhang, Li, and Bing 2023) $8B$ $22.71$
LLaVA-Next-Video (Zhang et al. 2024) $7B$ $22.90$
Qwen2–VL (Wang et al. 2024b) $7B$ $30.24$
X-CLIP-L/14-16F (Ni et al. 2022) $0.6B$ $16.26$ $33.74$ $49.89$
SigLIP–so400m (Zhai et al. 2023) (mean-pool) $0.9B$ $22.94$ $42.20$ $63.70$
+ DTW Alignment (Ours) $0.9B$ $23.27$ $45.66$ $67.37$
ActAlign (Ours) $0.9B$ $30.51$ $54.34$ $71.05$
Table 2: Zero-shot classification results on ActionAtlas (Salehi et al. 2024) under context-rich (T=0.2) prompting. ActAlign achieves state-of-the-art Top-1, Top-2, and Top-3 accuracy, outperforming all baselines and billion-parameter video–language models without any video–text supervision. These results highlight the effectiveness of structured sub-action alignment over flat representations such as mean-pooling and the open-set recognition capability of image-text models.
4.1 Experimental Setup
ActionAtlas Dataset
We construct SubActionAtlas upon the ActionAtlas benchmark, which— to the best of our knowledge—remains the most challenging dataset for fine-grained action recognition. It comprises 898 YouTube clips across 56 sports and 558 unique actions. Leveraging its rich action diversity and fine-grained complexity, we adopt ActionAtlas as the foundation for our zero-shot sequence-alignment evaluation. For each clip $V_{i}$, we retain its multiple-choice candidate set $\{c_{i,1},\dots,c_{i,M_{i}}\}$ but replace each high-level label with an LLM-generated sub-action sequence. This extension transforms ActionAtlas into a temporally grounded video–language alignment task, while preserving its original difficulty (human Top-1 accuracy: 61.64%).
Evaluation Metrics
We report Top-$k$ accuracy for $k\in\{1,2,3\}$:
[…]
where $\mathbb{I}$ is the indicator function and $\text{rank}_{i}(\hat{y}_{i})$ is the position of the ground‐truth label in the descending list of scores $\{\hat{\gamma}_{i,1},\dots,\hat{\gamma}_{i,M_{i}}\}$. This accounts for typical cases where fine-grained actions are semantically similar and alignment scores are closely clustered, allowing improvement to be captured within a narrowed candidate set.
Experimental Detail
We use SigLIP–so400m (patch size 14, $d=384$,  878M parameters). We apply a moving-average smoothing window of $w=30$ frames ($\approx$1s @30 fps) to reduce transient noise and emphasize consistent motion patterns. All experiments run on a single NVIDIA RTX A5000 GPU (25 GB).
Zero‐Shot Setup
Following the zero‐shot protocol, no example videos or sub-action sequence from these classes are used for any training or tuning. For each sample, we decompose each candidate action class $c_{m}$ into its sub-action sequence $S_{m}$ via LLM prompting, then align video frames $V_{i}$ against each sub-action sequence to compute the normalized alignment score $\hat{\gamma}_{i,m}$.
4.2 Experimental Result
Zero-Shot Baseline Comparisons
We first establish a random prediction baseline, followed by a zero-shot SigLIP baseline using mean-pooled frame embeddings, following ViFi-CLIP (Rasheed et al. 2023). Each video $V_{i}$ is represented by $\bar{\mathbf{z}}_{i}=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\mathbf{z}_{i}^{t}$, which is compared via cosine similarity to each class name embedding $\phi_{t}(c_{j})$.
We further compare against open-source video–language models, including entries from the ActionAtlas leaderboard and fine-tuned CLIP variants. Despite using no video–text supervision, our method improves over the SigLIP baseline by 7% Top-1 and 12% Top-2 accuracy, outperforming all baselines and billion-parameter video–language models with $\sim$8$\times$ fewer parameters.
As shown in Table 2, these gains are driven by our sub-action-based temporal alignment, which offers a discriminative representation than frame-level pooling alone.
Ablation Studies
We ablate each component of ActAlign on SubActionAtlas (Table 3), starting from a mean-pooled SigLIP baseline. Adding DTW alignment introduces temporal structure and yields consistent gains. Context augmentation—injecting domain context (e.g., “Sprint start” $\rightarrow$ “Sprint start in basketball”)—produces the largest boost by reducing semantic ambiguity. Signal smoothing offers a modest but complementary improvement by reducing frame-level noise and clarifying action boundaries.
Configuration Top-1 (%) Top-2 (%) Top-3 (%) SigLIP (Zhai et al. 2023) (mean-pool) $22.94$ $42.20$ $63.70$ + DTW Alignment $25.72$ $45.99$ $66.26$ + Context Augmentation $30.07$ $52.67$ $70.49$ + Signal Smoothing ($w=30$) $30.29$ $53.01$ $70.27$
Table 3: Ablation results under context-rich (T=1.0) prompting (see Table 1). DTW alignment introduces temporal matching, context augmentation reduces sub-action ambiguity, and signal smoothing mitigates frame-level noise.
We observe that the upper bound on performance is closely tied to the specificity and coherence of LLM-generated sub-action sequences.
Prompt Variations
We evaluate three prompt strategies for generating sub-action scripts using GPT-4o (see Table 1), keeping all other components fixed:
  • Short-Fixed ($T=1.0$): Prompts GPT-4o to generate exactly 10 terse (2–3 word) sub-actions per class using a fixed structure.
  • Context-Rich ($T=1.0$): Produces variable-length, context-rich sub-action scripts incorporating domain-specific cues (e.g., “Wrestler”, “Rider”) at standard temperature ($T=1.0$).
  • Context-Rich ($T=0.2$): Same as CR1.0 but with lower temperature to encourage more consistent and deterministic outputs.
Prompt Description Top-1 (%) Short‐Fixed 2-word, fixed 10 sub-actions $27.06$ Context‐Rich ($T=1.0$) context-rich, variable-length $30.29$ Context‐Rich ($T=0.2$) context-rich, variable-length (low $T$) 30.51
Table 4: Impact of prompt strategy. Context-rich prompting improves zero-shot classification performance by producing more specific and informative sub-actions.
Table 4 shows the performance of each strategy. The context-rich prompt with domain-specific context and low temperature ($T=0.2$) achieve the highest accuracy. In contrast, short-fixed prompts—lacking sufficient semantic specificity—perform worst. These results highlight that reducing ambiguity in sub-action descriptions directly improves alignment quality and classification performance.
Sub-action Sequence Examples
Table 5 shows LLM-generated subaction sequences for two Figure skating tactics under our context-rich ($T=0.2$) prompt. The scripts highlight ordered, salient steps enabling precise temporal matching.
Tactic-Subaction Script Biellmann Spin 1. Begins upright spin on one foot with arms extended and free leg behind, 2. Gradually pulls free leg upward behind the back using both hands, 3. Raises the free leg above head level while arching the back dramatically, 4. Grasps the blade of the free skate with both hands overhead, 5. Extends spinning leg vertically while maintaining centered spin on skating foot, 6. Maintains high-speed rotation with body in extreme vertical split position Flying Camel Spin 1. Skater glides forward with arms extended and knees bent in preparation, 2. Performs a powerful jump off the toe pick while swinging free leg upward, 3. Rotates mid-air with body extended horizontally like a ’T’ shape, 4. Lands on one foot directly into a camel spin position with torso parallel to ice, 5. Extends free leg backward and arms outward while spinning on the skating leg, 6. Maintains fast, centered rotation in the horizontal camel position
Table 5: LLM-generated sub-action scripts for figure skating tactics. Shown for the Biellmann Spin and Flying Camel Spin examples in Figure 4, these sequences are generated using context-rich prompting ($T=0.2$) and provide semantically detailed, temporally ordered steps for alignment in our zero-shot framework.
Alignment Heatmaps and Paths
Figure 4 visualizes the cosine similarity matrix and DTW path for a correctly classified action (right) and an incorrect candidate (left). The correct sequence yields high-similarity regions with a monotonic path. In contrast, the incorrect script shows sparse similarity regions, with the DTW path forced to follow the single most similar alignment trace.
Video Example: Performing flying camel spin in figure skating
Biellmann spin $\times$     (Incorrect)
Flying camel spin $\checkmark$     (Correct)
Figure 4: Comparison of similarity heatmaps and DTW alignment paths for a correct classification (right) versus an incorrect prediction (middle). The correct class exhibits clearer segmentation and higher alignment quality. The sub-action scripts are provided in Table 5.
Figure 5: t-SNE visualization of sub-action embeddings without (left) and with (right) context augmentation under Context-Rich prompting ($T=0.2$). Colors indicate different sport domains.
Error Analysis
We identify two primary failure modes in our framework:
  • Ambiguous sub-actions: When LLM-generated scripts include vague steps (e.g., “move to position”), the similarity matrix becomes sparse, limiting DTW’s ability to discriminate between classes. As shown in Figure 6, context-rich prompts yield high-similarity regions than short-fixed ones. Appendix Table 2 highlights the qualitative difference in sub-action specificity.
  • Global alignment bias: DTW enforces full-sequence alignment, which may fail when the action begins mid-clip or exhibits temporal shifts. Without a local alignment mechanism, early or trailing sub-actions introduce noise. Figure 4 illustrates a temporal shift case that does not harm classification in this instance, but could lead to errors when candidate classes are highly similar.
These findings highlight the need for precise sub-action generation and motivate improvements in action prototype design.
(a) Video Snapshot: Representative frame showing the execution of Varial kickflip underflip body varial in Skateboarding.
(b) Short‐Fixed (incorrect, $\hat{\gamma}=0.35$).
(c) Context‐Rich ($T=0.2$, correct, $\hat{\gamma}=0.81$).
Figure 6: Failure case due to vague sub-actions for the action Varial kickflip underflip body varial. (6(b)) The Short-Fixed prompt yields a noisy similarity map and incorrect prediction. (6(c)) The Context-Rich prompt produces clearer alignment and correct classification. See Appendix Table 2 for scripts.
Sub-action Embedding Structure
Figure 5 shows 2D t-SNE projections of sub-action embeddings with and without context augmentation under context-rich prompts ($T=0.2$). Adding domain-specific cues ({sub-action} in {sport name}) results in tighter, more coherent clusters—indicating better semantic structure and separation.
Effect of Signal Smoothing
Figure 7 compares similarity matrices before and after applying a moving-average filter. Without smoothing, rapid scene changes introduce noise and scattered peaks. Smoothing yields cleaner similarity surfaces with clearer action boundaries.
(a) Without smoothing
(b) With smoothing ($w=30$)
Figure 7: Similarity matrices before and after applying a moving-average filter ($w=30$). Signal smoothing reduces high-frequency noise and enhances transition between sub-actions.
5 Conclusion
We show that contrastive image–text models, even when used with simple mean-pooling, establish a surprisingly strong baseline for open-set fine-grained video classification. To fully leverage this capability in extremely fine-grained settings, we propose ActAlign, a novel zero-shot framework that revisits the classic Dynamic Time Warping algorithm to cast video classification as a sequence-matching problem. By aligning video frames with LLM-generated sub-action scripts, ActAlign introduces temporal structure into contrastive models without requiring any video–text training or fine-tuning. Evaluated on the highly challenging ActionAtlas benchmark, our method achieves state-of-the-art performance, outperforming both CLIP-style baselines and billion-parameter video–language models. These results underscore the value of structured language priors and classical alignment in unlocking the open-set recognition potential of vision–language models for fine-grained video classification.
Future Work.
The effectiveness of ActAlign remains limited by the quality of LLM-generated sub-action scripts. Future work includes exploring improved prompting strategies and local alignment techniques. Given its general design, ActAlign can be integrated with any vision–language model to advance zero-shot fine-grained video understanding across other domains.